ECE 3 1. Simplify the following expression a. ;33 = b. 2; +3; = c. (51) = d. ;; = f. 3 (2j? -5j)= g. 6xj? +2 yj* + 4.xj +
Answer #1

Note: As per Chegg policy we have to answer first four parts only.

By using j^2=-1 we get the solution.

23 (ه ن 32 - 33 والان) = We know j² = -1 الان : نه کارت 1 1xj = زة +زء (ط نذ 5 + == [۱- = ذ :C از3- 2- = 3 + 2

روزه) و وز 53 = xjx 5ڑا = xj(ن) کا = لا = ذ :] x - xj 125 = (زی) 125 x ( - = 125 تن را گ(ن) = = [1- ن :] 1- = ( = ان - - ن

نبهان = ز (ن) ۔ دا - لان :C ن = xj راه = وذ (5 - نحازة ( از 15 ۔ زنی - [ - -ان :C 15 + ز = زه - 15 =

- 63x 9) 6x j® +2y;+ + 4x3 +54 58 = 6z j&j + 24 ()+423 +5°(3%) tij?--1] =-bxjtzy (-1) + 4xj+54 (-1) = - 6zj +2y + 423 +54 = 7